# Talha Bozkus and Urbashi Mitra

### v2

@@ Abstract
$Q$-learning is a widely used reinforcement learning (RL) algorithm for optimizing wireless networks, but faces challenges with large state-spaces. Recently proposed multi-environment mixed $Q$-learning (MEMQ) algorithm addresses these challenges by employing multiple $Q$-learning algorithms across multiple synthetically generated, distinct but structurally related environments, so-called digital cousins. In this paper, we propose a novel multi-agent MEMQ (M-MEMQ) for cooperative decentralized wireless networks with multiple networked transmitters (TXs) and base stations (BSs). TXs do not have access to global information (joint state and actions). The new concept of coordinated and uncoordinated states is introduced. In uncoordinated states, TXs act independently to minimize their individual costs and update local $Q$-functions. In coordinated states, TXs use a Bayesian approach to estimate the joint state and update the joint $Q$-functions. The cost of information-sharing scales linearly with the number of TXs and is independent of the joint state-action space size. Several theoretical guarantees, including deterministic and probabilistic convergence, bounds on estimation error variance, and the probability of misdetecting the joint states, are given. Numerical simulations show that M-MEMQ outperforms several decentralized and centralized training with decentralized execution (CTDE) multi-agent RL algorithms by achieving 55% lower average policy error (APE), 35% faster convergence, 50% reduced runtime complexity, and 45% less sample complexity. Furthermore, M-MEMQ achieves comparable APE with significantly lower complexity than centralized methods. Simulations validate the theoretical analyses.
@@ I Introduction
Reinforcement learning (RL) is a powerful framework for optimizing next-generation wireless networks that can be modeled as Markov Decision Processes (MDPs) [1, 2, 3]. When the MDP dynamics, such as transition probabilities and cost functions, are unknown, model-free RL algorithms such as $Q$-learning can be employed. $Q$-learning is a simple, tabular, and model-free RL algorithm and works with discrete and modest-sized state-action spaces [4]. However, it faces several challenges, including slow convergence, high sample complexity, high estimation bias, and variance, and training instability across large state-action spaces [5, 6].
@@
We have proposed multi-environment mixed $Q$-learning (MEMQ) to tackle these issues by running multiple $Q$-learning algorithms across multiple, distinct yet structurally related synthetic environments – digital cousins and fusing their $Q$-functions into a single ensemble $Q$-function estimate using adaptive weighting [5, 6, 7, 8, 9]. MEMQ has three distinct features compared to the existing $Q$-learning algorithms. First, MEMQ constructs multiple, synthetic, distinct yet structurally related environments during training. These environments have been shown to help accelerate the exploration of the state-space [5]. Second, MEMQ trains multiple $Q$-learning agents on these environments and combines their outputs into an ensemble $Q$-function to achieve more robust $Q$-functions with lower variance [5]. Third, MEMQ uses a mixed-learning strategy – it combines real-time interactions with synthetic samples collected from the synthetic environments for faster training with improved sample complexity [5, 6].
@@
MEMQ was originally designed for single-agent centralized MISO and MIMO wireless networks to solve real-world challenges such as resource allocation and interference minimization with lower complexity. However, it is not directly applicable to multi-agent and decentralized wireless networks. To this end, multi-agent RL (MARL) theory can be used to extend single-agent MEMQ to multi-agent networks. MARL has been effectively applied to different wireless network optimization problems, including resource and power allocation, channel access, beamforming, spectrum management, and multi-access edge computing [10, 11, 12, 13, 14, 15, 16], and can be categorized along several dimensions. Different training strategies depend on how agents are trained and execute their decisions and include fully centralized [5, 17], centralized training decentralized execution (CTDE) [18, 19], decentralized with networked agents (DNA) [20, 21] and fully decentralized [22, 23]. If agents have complete visibility of the global state, it is fully observable; otherwise partially observable [24]. Different communication strategies may vary in terms of which agents communicate, when and how to establish communication links, what information to share, and how to combine the received messages [25]. Different coordination strategies depend on whether agents collaborate to achieve a common goal and include fully cooperative [26], fully competitive [27], or mixed [28]. If agents share the same observation and action space and have a common reward function, then they are homogeneous [29]; otherwise heterogeneous [30]. Other categorizations include value-based vs. policy-based methods, tabular vs. deep neural network approaches, online vs. offline vs. hybrid methods, and methods for discrete vs. continuous state-action spaces.
@@
MARL face challenges such as partial observability, scalability, effective exploration and security concerns, non-stationary, and credit assignment [31]. Our proposed M-MEMQ algorithm can alleviate these challenges effectively. For example, while global information is not available to agents, they use MAP estimation based on their local observations to estimate the joint state. On the other hand, M-MEMQ offers a good balance between policy accuracy, runtime complexity, and communication costs across an increasing number of agents and larger state-action space, making it scalable. M-MEMQ also accelerates the exploration of the joint state-space by combining real-time interactions with synthetic samples collected from synthetic environments in real-time. The problem of non-stationary is alleviated through the leader agent, which uses a centralized $Q$-function update by aggregating information from individual agents to maintain a global system view. Finally, the leader agent updates the global $Q$-functions and determines the joint state and action, considering each agent’s confidence in their joint state estimates, which prioritizes more confident agents by assigning them greater weight (i.e., credit) for their contributions.
@@
Our approach herein is a form of generative learning, but differs from traditional generative methods such as generative adversarial networks [32, 33, 34]. We construct synthetic environments – synthetic simulators that RL agents can continuously interact with to generate synthetic training data alongside real data during training. Our approach is computationally efficient to implement, interpretable, and has strong theoretical guarantees. The synthetically generated data herein follow different, but related distribution to the real-time data and are augmented with real-time data to accelerate exploration [5, 6]. We observe that due to our exemplar application, wireless communication networks, our distinct agents actually experience different environments and dynamics in contrast to many MARL works [10, 11, 12].
@@
The main contributions of this paper are as follows: (i) We develop the first multi-agent MEMQ algorithm (M-MEMQ) for decentralized wireless networks with networked agents. (ii) In coordinated states, agents estimate the joint state using a Bayesian MAP estimation based on local observations in a decentralized way to minimize the cost of information exchange. In the uncoordinated states, agents act independently. (iii) We provide the first theoretical guarantees for multi-agent MEMQ algorithms, including deterministic and probabilistic convergence, bounds on estimation error variance, and the probability of misdetecting the joint states. (iv) We evaluate our algorithm on a wireless network with multiple mobile transmitters (TX) and stationary base stations (BS) to find the optimal TX-BS association to minimize the joint wireless cost. Our algorithm outperforms several decentralized and CTDE-based MARL algorithms by achieving 55% lower average policy error (APE), 35% faster convergence rate, 50% less computational runtime complexity and %45 less sample complexity. Our algorithm achieves comparable APE with significantly lower complexity than centralized methods. In addition, our algorithm exhibits high robustness to changes in network parameters. Simulations verify theoretical analyses.
@@ II-A Markov Decision Processes
A multi-agent MDP generalizes a single-agent MDP to $N$ agents ($i$ = $1,2,...,N$), each with its own states and actions. The joint state at time $t$ is $\bar{s}_{t}=(s_{1,t},\ldots,s_{N,t})$ and the joint action is $\bar{a}_{t}=(a_{1,t},\ldots,a_{N,t})$. The joint state and action spaces are obtained as $\mathcal{S}=\otimes_{i=1}^{N}\mathcal{S}_{i}$ and $\mathcal{A}=\otimes_{i=1}^{N}\mathcal{A}_{i}$, where $\otimes$ is the Kronecker product. The transition from joint state $\bar{s}_{t}$ to $\bar{s}^{\prime}_{t}$ under joint action $\bar{a}_{t}$ occurs with probability $p(\bar{s}_{t},\bar{a}_{t},\bar{s}^{\prime}_{t})$, which is stored in the joint probability transition tensor $\mathbf{\bar{P}}$, and incurs joint cost $c_{i}(\bar{s}_{t},\bar{a}_{t})$ for agent $i$. The optimization problem of the multi (joint) agent is as follows:
@@ II-B $Q$-Learning
where $\alpha\in(0,1)$ is the learning rate. Different exploration strategies, such as $\epsilon$-greedy, softmax-exploration, or upper-confidence bound action selection, can be used to handle the exploration-exploitation trade-off to ensure sufficient sampling of different state-action pairs [4]. The agent interacts with the environment and collects samples $\{s_{i},a_{i},s^{\prime}_{i},c_{i}\}$ to update $Q$-functions using (3). $Q$-functions converge to their optimal values with probability one, given certain conditions [35]. The optimal policy is $\pi_{i}^{*}(s_{i})=\operatorname*{argmin}_{a_{i}\in\mathcal{A}_{i}}Q_{i}^{*}(s_{i},a_{i})$, and the relationship between the optimal value function and optimal $Q$-functions is $v_{i}^{*}(s_{i})=\min_{a_{i}\in\mathcal{A}_{i}}Q_{i}^{*}(s_{i},a_{i})$. $Q$-learning is an off-policy algorithm because it optimizes a target policy (the optimal policy being learned) while using a different policy, such as $\epsilon$-greedy, for exploration and interaction with the environment. In addition, $Q$-learning effectively works with small to moderately large discrete state-action spaces.
@@ II-C Single-agent Multi-environment Mixed $Q$-Learning
We herein briefly review the work of [5]. The agent $i$ first estimates the PTT of the real environment in real-time by sampling. Let $\mathbf{\hat{P}}_{i}$ denote the estimated PTT. Simultaneously, it constructs the PTT of synthetic environments as powers of $\mathbf{\hat{P}}_{i}$, with $\mathbf{\hat{P}}^{n}_{i}$ representing the estimated PTT of the $n$-th synthetic environment  [^1] There are alternative ways to construct the PTT of synthetic environments such as employing graph symmetrization methods [6].. The agent runs $K_{i}$ independent $Q$-learning algorithms across these environments, generating $Q$-functions $Q^{(n)}_{i}$ for each $\text{Env}_{n}$. These $Q$-functions are combined into an ensemble $Q$-function, $Q^{e}_{i}$, using the update rule:
@@
where the weights of environments $w_{i}^{(n)}$ is determined using some adaptive weighting mechanism, and $u$ is the update ratio parameter [5]. It has been shown that $Q^{e}_{i}$ converges to the optimal $Q$-functions in mean-square (i.e. $Q^{e}_{i}\xrightarrow{m.s.}Q^{*}$ for all $i$) faster than the individual $Q$-functions $Q^{(n)}_{i}$ given a particular structure on update ratio $u$ [5].
Herein, $K_{i}$ and $n$ are hyperparameters. While $K_{i}$ is chosen proportional to the state-action space size, selecting $n$ is more challenging. For $K_{i}=4$, there are many possible sets of orders (1,2,3,4 vs 1,2,3,5 vs 2,3,6,7 or similar). The goal is to use the fewest synthetic environments that provide the most diverse multi-time-scale information. To this end, we employ the coverage-coefficient environment selection algorithm that we developed for the single-agent case [36, 37], which selects the synthetic environments based on their sample coverage of the real environment.
@@ III Decentralized Wireless Network with Multiple Networked Agents
While our approach is general and has wide applications [5, 6], to make our approach and analysis more accessible, we first describe a decentralized wireless network framework with multiple agents, serving as a representative example of real-world network complexities and challenges.
We consider a grid environment of size $L\times L$ with cells of size $\Delta_{L}\times\Delta_{L}$. There are $N_{T}$ mobile transmitters (TXs) and $N_{B}$ stationary base stations (BSs), where each TX can move randomly in four directions (U, R, D, L) or stay stationary, all with equal probability. TXs always remain on grid points. Each BSj has a circular coverage area with radius $r_{j}$; a TX can connect to BS only on or within this area. Fig.1 shows an example wireless network with 3 TXs and 2 BSs, where red squares are BSs, green triangles are TXs, black solid circles show coverage areas, purple solid arrows represent a particular TX-BS association, black dotted arrows represent the communications between TXs, and TX1 is assumed to be the leader TX that controls the coordination among agents (i.e. TXs communicate with only TX1 when necessary). Herein, the leader TX is defined as the transmitter with the largest coverage area (radius $r_{j}$), as it can communicate with and broadcast to the largest portion of the network, ensuring better connectivity to base stations and other TXs. We underline that many hierarchical sensor and wireless networks have leader transmitters or nodes [38]. Each TX can continuously measure (sense) the aggregated received signal strength (ARSS) due to all other TXs. This is practical, as most wireless devices can easily obtain ARSS measurements, unlike CSI information, which requires specialized hardware [39]. We now describe the RL framework for the given wireless network.

### v3

@@ Abstract
$Q$-learning is a widely used reinforcement learning (RL) algorithm for optimizing wireless networks, but faces challenges with large state-spaces. Recently proposed multi-environment mixed $Q$-learning (MEMQ) algorithm addresses these challenges by employing multiple $Q$-learning algorithms across multiple synthetically generated, distinct but structurally related environments, so-called digital cousins. In this paper, we propose a novel multi-agent MEMQ (M-MEMQ) for cooperative decentralized wireless networks with multiple networked transmitters (TXs) and base stations (BSs). TXs do not have access to global information (joint state and actions). The new concept of coordinated and uncoordinated states is introduced. In uncoordinated states, TXs act independently to minimize their individual costs and update local $Q$-functions. In coordinated states, TXs use a Bayesian approach to estimate the joint state and update the joint $Q$-functions. The cost of information-sharing scales linearly with the number of TXs and is independent of the joint state-action space size. Several theoretical guarantees, including deterministic and probabilistic convergence, bounds on estimation error variance, and the probability of misdetecting the joint states, are given. Numerical simulations show that M-MEMQ outperforms several decentralized and centralized training with decentralized execution (CTDE) multi-agent RL algorithms by achieving 60% lower average policy error (APE), 40% faster convergence, 45% reduced runtime complexity, and 40% less sample complexity. Furthermore, M-MEMQ achieves comparable APE with significantly lower complexity than centralized methods. Simulations validate the theoretical analyses.
@@ I Introduction
Reinforcement learning (RL) is a powerful framework for optimizing next-generation wireless networks that can be modeled as Markov Decision Processes (MDPs) [1, 2]. When the MDP dynamics, such as transition probabilities and cost functions, are unknown, model-free RL algorithms such as $Q$-learning can be employed. $Q$-learning is a simple, tabular, and model-free RL algorithm and works with discrete and modest-sized state-action spaces [3]. However, it faces several challenges, including slow convergence, high sample complexity, high estimation bias, and variance, and training instability across large state-action spaces [4, 5].
@@
We have proposed multi-environment mixed $Q$-learning (MEMQ) to tackle these issues by running multiple $Q$-learning algorithms across multiple, distinct yet structurally related synthetic environments – digital cousins and fusing their $Q$-functions into a single ensemble $Q$-function estimate using adaptive weighting [4, 5, 6]. MEMQ has three distinct features compared to the existing $Q$-learning algorithms. First, MEMQ constructs multiple, synthetic, distinct yet structurally related environments during training. These environments have been shown to help accelerate the exploration of the state-space [4]. Second, MEMQ trains multiple $Q$-learning agents on these environments and combines their outputs into an ensemble $Q$-function to achieve more robust $Q$-functions with lower variance [4]. Third, MEMQ uses a mixed-learning strategy – it combines real-time interactions with synthetic samples collected from the synthetic environments for faster training with improved sample complexity [4, 5].
@@
MEMQ was originally designed for single-agent centralized MISO and MIMO wireless networks to solve real-world challenges such as resource allocation and interference minimization with lower complexity. However, it is not directly applicable to multi-agent and decentralized wireless networks. To this end, multi-agent RL (MARL) theory can be used to extend single-agent MEMQ to multi-agent networks. MARL has been effectively applied to different wireless network optimization problems, including resource and power allocation, channel access, beamforming, spectrum management, and multi-access edge computing [7, 8, 9, 10], and can be categorized along several dimensions. Different training strategies depend on how agents are trained and execute their decisions and include fully centralized [4, 11], centralized training decentralized execution (CTDE) [12], decentralized with networked agents (DNA) [13, 14] and fully decentralized [15, 16]. If agents have complete visibility of the global state, it is fully observable; otherwise partially observable [17]. Different communication strategies may vary in terms of which agents communicate, when and how to establish communication links, what information to share, and how to combine the received messages [18]. Different coordination strategies depend on whether agents collaborate to achieve a common goal and include fully cooperative [19], fully competitive [20], or mixed [21]. If agents share the same observation and action space and have a common reward function, then they are homogeneous [22]; otherwise heterogeneous [23]. Other categorizations include value-based vs. policy-based methods, tabular vs. deep neural network approaches, online vs. offline vs. hybrid methods, and methods for discrete vs. continuous state-action spaces.
@@
MARL algorithms face challenges such as partial observability, scalability, effective exploration and security concerns, non-stationary, and credit assignment [24]. Our proposed M-MEMQ algorithm can alleviate these challenges effectively. For example, while global information is not available to agents, they use MAP estimation based on their local observations to estimate the joint state. On the other hand, M-MEMQ offers a good balance between policy accuracy, runtime complexity, and communication costs across an increasing number of agents and larger state-action space, making it scalable. M-MEMQ also accelerates the exploration of the joint state-space by combining real-time interactions with synthetic samples collected from synthetic environments in real-time. The problem of non-stationary is alleviated through the leader agent, which uses a centralized $Q$-function update by aggregating information from individual agents to maintain a global system view. Finally, the leader agent updates the global $Q$-functions and determines the joint state and action, considering each agent’s confidence in their joint state estimates, which prioritizes more confident agents by assigning them greater weight (i.e., credit) for their contributions.
Our multi-agent MEMQ algorithm introduces substantial improvements over the single-agent MEMQ framework [4, 5]. We employ a minimal-overhead communication protocol where agents share only local Q-values and costs, ensuring scalability linear in the number of agents rather than the joint state–action space [25]. To overcome partial observability, we use Bayesian joint-state estimation from local observations without requiring state information exchanges [25]. We systematically evaluate and optimize different digital cousin architectures (shared, common, independent) for decentralized environments. Our algorithm supports real-time adaptive coordination, dynamically switching between coordinated and uncoordinated modes using ARSS measurements instead of fixed assumptions [9, 26]. Finally, we design MEMQ-specific Q-function update rules tailored to the multi-environment setting, which differ fundamentally from prior MARL. We also introduce new theoretical contributions that are mostly missing in the partially decentralized MARL literature [27, 28]. We establish both probabilistic and deterministic convergence results and upper bounds on the variance of estimation errors for the joint $Q$-functions, and provide sample complexity guarantees for achieving accurate policies. We further analyze the criterion for distinguishing coordinated from uncoordinated states and quantify the probability of misdetection under noisy ARSS, whereas prior works treat this selection as fixed or heuristic without theoretical justifications [29, 9].
@@
Our approach herein is a form of generative learning, but differs from traditional generative methods such as generative adversarial networks [30]. We construct synthetic environments – synthetic simulators that RL agents can continuously interact with to generate synthetic training data alongside real data during training. Our approach is computationally efficient to implement, interpretable, and has strong theoretical guarantees. The synthetically generated data herein follow different, but related distribution to the real-time data and are augmented with real-time data to accelerate exploration [4, 5]. We observe that due to our exemplar application, wireless communication networks, our distinct agents actually experience different environments and dynamics in contrast to many MARL works [7, 8].
@@
The main contributions of this paper are as follows: (i) We develop the first multi-agent MEMQ algorithm (M-MEMQ) for decentralized wireless networks with networked agents. (ii) In coordinated states, agents estimate the joint state using a Bayesian MAP estimation based on local observations in a decentralized way to minimize the cost of information exchange. In the uncoordinated states, agents act independently. (iii) We provide the first theoretical guarantees for multi-agent MEMQ algorithms, including deterministic and probabilistic convergence, bounds on estimation error variance, and the probability of misdetecting the joint states. (iv) We evaluate our algorithm on a wireless network with multiple mobile transmitters (TX) and stationary base stations (BS) to find the optimal TX-BS association to minimize the joint wireless cost. Our algorithm outperforms several decentralized and CTDE-based MARL algorithms by achieving 60% lower average policy error (APE), 40% faster convergence rate, 45% less computational runtime complexity and 40% less sample complexity. Our algorithm achieves comparable APE with significantly lower complexity than centralized methods. In addition, our algorithm exhibits high robustness to changes in network parameters. Simulations verify theoretical analyses.
@@ II-A Markov Decision Processes
[…]
@@ II-B $Q$-Learning
where $\alpha\in(0,1)$ is the learning rate. Different exploration strategies, such as epsilon-greedy, softmax-exploration, or upper-confidence bound action selection, can be used to handle the exploration-exploitation trade-off to ensure sufficient sampling of different state-action pairs [3]. The agent interacts with the environment and collects samples $\{s_{i},a_{i},s^{\prime}_{i},c_{i}\}$ to update $Q$-functions using (3). $Q$-functions converge to their optimal values with probability one, given certain conditions [31]. The optimal policy is $\pi_{i}^{*}(s_{i})=\operatorname*{argmin}_{a_{i}\in\mathcal{A}_{i}}Q_{i}^{*}(s_{i},a_{i})$, and the relationship between the optimal value function and optimal $Q$-functions is $v_{i}^{*}(s_{i})=\min_{a_{i}\in\mathcal{A}_{i}}Q_{i}^{*}(s_{i},a_{i})$. $Q$-learning is an off-policy algorithm because it optimizes a target policy (the optimal policy being learned) while using a different policy, such as epsilon-greedy, for exploration and interaction with the environment. In addition, $Q$-learning effectively works with small to moderately large discrete state-action spaces.
@@ II-C Single-agent Multi-environment Mixed $Q$-Learning
We herein briefly review the work of [4]. The agent $i$ first estimates the PTT of the real environment in real-time via sample averaging from empirical state transitions collected under the $\varepsilon$-greedy exploration policy [32]. This approach produces Maximum Likelihood Estimate (MLE) for the PTTs unless additional structure exists [32, 33]. Let $\mathbf{\hat{P}}_{i}$ denote the estimated PTT. This estimate $\hat{\mathbf{P}}_{i}$ is then used to construct synthetic transition kernels for the digital cousins. Specifically, the $n$-step transition kernel of the $n^{\text{th}}$-order digital cousin is computed as $\hat{\mathbf{P}}_{i}^{n}$.  [^1] There are alternative ways to construct the PTT of synthetic environments such as employing graph symmetrization methods [5].. The agent runs $K_{i}$ independent $Q$-learning algorithms across these environments, generating $Q$-functions $Q^{(n)}_{i}$ for each $\text{Env}_{n}$. These $Q$-functions are combined into an ensemble $Q$-function, $Q^{e}_{i}$, using the update rule:
@@
where the weights of environments $w_{i}^{(n)}$ is determined using some adaptive weighting mechanism, and $u$ is the update ratio parameter [4]. It has been shown that $Q^{e}_{i}$ converges to the optimal $Q$-functions in mean-square (i.e. $Q^{e}_{i}\xrightarrow{m.s.}Q^{*}$ for all $i$) faster than the individual $Q$-functions $Q^{(n)}_{i}$ given a particular structure on update ratio $u$ [4]. We emphasize that neither the synthetic $Q$-functions $Q_{i}^{(n)}$ nor their weighted combinations $\sum_{n=1}^{K_{i}}w_{i}^{(n)}Q_{i}^{(n)}$ generally converge to the optimal $Q$-function due to different underlying PTTs. Instead, the ensemble $Q$-function $Q_{i}^{e}$, which is obtained iteratively via update rule (4), does converge to the optimal $Q$-function [4].
Herein, $K_{i}$ and $n$ are hyperparameters. While $K_{i}$ is chosen proportional to the state-action space size, selecting $n$ is more challenging. While the parameter $n$ can range from $1$ to $K_{i}$ as in (4), in practice, only a subset of these values is selected, since not all $K$ environments contribute useful samples. For $K_{i}=4$, there are many possible sets of orders (1,2,3,4 vs 1,2,3,5 vs 2,3,6,7 or similar). The goal is to use the fewest synthetic environments that provide the most diverse multi-time-scale information. To this end, we employ the coverage-coefficient environment selection algorithm that we developed for the single-agent case [34, 35], which selects the synthetic environments based on their sample coverage of the real environment. We also underline that, regardless of $K_{i}$, samples from synthetic environments cannot contain more information than real samples; hence, the real environment is always included in the MEMQ algorithm.
@@ III Decentralized Wireless Network with Multiple Networked Agents
While our approach is broadly applicable to domains such as wireless network optimization, multi-agent games, and random network graph optimization (as demonstrated in the single-agent case [4, 5]), in this work we concentrate on decentralized wireless networks. This setting provides a representative example of real-world network complexities and challenges, while the framework remains readily adaptable to games, network graphs, and other applications.
We consider a grid environment of size $L\times L$ with cells of size $\Delta_{L}\times\Delta_{L}$. There are $N_{T}$ mobile transmitters (TXs) and $N_{B}$ stationary base stations (BSs), where each TX can move randomly in four directions (U, R, D, L) or stay stationary, all with equal probability. TXs always remain on grid points. Each BSj has a circular coverage area with radius $r_{j}$; a TX can connect to BS only on or within this area. Fig.1 shows an example wireless network with 3 TXs and 2 BSs, where red squares are BSs, green triangles are TXs, black solid circles show coverage areas, purple solid arrows represent a particular TX-BS association, black dotted arrows represent the communications between TXs, and TX1 is assumed to be the leader TX that controls the coordination among agents (i.e. TXs communicate with only TX1 when necessary). Herein, the leader TX is defined as the transmitter with the largest coverage area (radius $r_{j}$), as it can communicate with and broadcast to the largest portion of the network, ensuring better connectivity to base stations and other TXs. We underline that many hierarchical sensor and wireless networks have leader transmitters or nodes [36]. We emphasize that although leader selection introduces minor overhead, the leader need not be fixed and can be reassigned across simulations; as long as a central agent with sufficiently large coverage exists, our analysis remains valid. Each TX can continuously measure (sense) the aggregated received signal strength (ARSS) due to all other TXs. This is practical, as most wireless devices can easily obtain ARSS measurements, unlike CSI information, which requires specialized hardware [37]. We now describe the RL framework for the given wireless network.